pH of 1L of aqueous solution of HCL is 1,then what amount of water to be added to change pH to 2?

Grade:11

1 Answers

Pankaj
9 years ago
Molarity of HCl at 1 pH
pH = -log[H+]
1 = -log10[H+]
or [H+] = 0.1
Similarly
Molarity of HCl at pH 2 = 0.02
Now
M1V1 = M2V2
or
1 x 0.1 = 0.01 x V2
or
V2 = 10
So, amouth of water atted = V2-V1 = 10-1 – 9 L